(x+y-4)dx-(3x-y-4)dy=0 when x=4 and y=7.


1
Expert's answer
2021-04-15T07:27:05-0400

(x+y-4)dx-(3x-y-4)dy=0 

"\frac{dy}{dx} = \frac{x+y-4}{3x-y-4}"

Let x = u+h and. y = v+k

"\frac{x+y-4}{3x-y-4}" = "\frac{u+h+v+k-4}{3u+3h-v-k-4}=\frac{u+v+h+k-4}{3u-v+3h-k-4}"

To make it homogeneous put

h+k-4=0 and 3h-k-4=0

Adding we get 4h = 8=> h = 2

So k = 2

So x = u+2 and y = v+2

dx = du and dy = dv

So "\frac{dy}{dx} = \frac{dv}{du}"

Now the given differential equation transforms to

"\frac{dv}{du} = \frac{u+v}{3u-v}"

Let v = tu

"\frac{dv}{du} = t + u \frac{dt}{du}"

=> "t + u \frac{dt}{du} =" "\frac{u+tu}{3u-tu}"

=> "u \frac{dt}{du} =" "\frac{1+t}{3-t}-t"

=> "u \frac{dt}{du} =" "\frac{1-2t+t²}{3-t}"

=> "\frac{3-t}{(t-1)²} dt = \frac{du}{u}"

=> "\frac{2-(t-1)}{(t-1)²} dt = \frac{du}{u}"

=> "\frac{2}{(t-1)²} dt -\frac{1}{(t-1)} dt=\frac{du}{u}"

=> "\int\frac{2}{(t-1)²} dt -\int\frac{1}{(t-1)} dt=\int\frac{du}{u}"

=> "-\frac{2}{t-1} - ln|t-1|=ln|u| - lnC"

=> "-\frac{2}{t-1} = ln|\frac{u(t-1)}{C}|"

=> "-\frac{2u}{v-u} = ln|\frac{(v-u)}{C}|"

=> "-\frac{2(x-2)}{y-x} = ln|\frac{(y-x)}{C}|"

=> "\frac{(y-x)}{C}=" "e^{-\frac{2(x-2)}{y-x} }"

=> y = x + "Ce^{-\frac{2(x-2)}{y-x} }" , This is the general solution.

Now y = 7 when x = 4

So 7 = 4 + "Ce^{-\frac{2(4-2)}{7-4} }"

=> C = 3"e^{\frac{4}{3} }"

So the solution is

y = x "+e^{\frac{4}{3}}.e^{-\frac{2(x-2)}{y-x} }"